OpenStudy (anonymous):

Find the exact value of each real number y if it exists. y=sin^-1 √2/2

OpenStudy (joannablackwelder):

You can rewrite as siny =sqrt2/2. Since siny=opp/hyp you can draw a triangle with angle y.

OpenStudy (joannablackwelder):

I recognize this triangle as ratios of a 45-45-90 triangle since they have a ratio of x,x, xsqrt(2).

OpenStudy (joannablackwelder):

So, y is 45 deg.

OpenStudy (anonymous):

thank you so much. you are so helpful. :*)

OpenStudy (joannablackwelder):

No worries. Glad I could help. :)